# Structural response of neutron stars to rapid rotation and its impact on the braking index

### v1

@@ V Conclusion
In this paper, we have investigated the effect of stellar rotation on the braking index and its dependence on the EOS. We have generated a large number of EOS by employing a semi-agnostic model and constrained them using the latest astrophysical observations from the gravitational and radio wave bands within a Bayesian framework. We find our constrained parameter space is also consistent with the observational findings of neutron star’s mass and radius from X-ray bands. The constrained nuclear EOSs have been used to study the effect on the macroscopic quantities of rapidly rotating neutron stars like mass, radius, moment of inertia and the ratio of polar to equatorial radius. Our analysis shows neutron stars with stiffer EOS experience larger deformations in the shape at a given frequency, in comparison to the softer EOS. Consequently, neutron stars with stiffer EOS have smaller Kepler frequencies.
@@
Next, we study the effect of rotation on the braking index of neutron stars. We present the first calculation of the braking index accounting for the rotational effects in the presence of energy loss by magnetic dipole radiation and GWs due to non-axisymmetric deformations and r-mode oscillations. We have shown that the rotation of the star can impact its braking index significantly. The deviation of the braking index from 3 to smaller values in the case of magnetic dipole radiation is due to the rotational effects. This rotational effect is prominent only when the spin frequency $\gtrsim 200$ Hz. However, the low braking index in normal pulsars has been explained through various other mechanisms, such as, the evolution of magnetic inclination angle, which has not been considered here. Similarly, we have shown if the neutron stars spin-down is purely due to GW radiation from non-asymmetric deformations or r-mode oscillations, the effect of rapid rotation leads to the significant departure of braking index from their classical value of 5 and 7 respectively.
@@
We further explore the dependence of the braking index on the rotation frequency for various physically motivated combinations of surface magnetic field strength, ellipticity of the star and the amplitude of r-mode oscillations. We find if there is r-mode instability generated in a typical millisecond pulsar with $\alpha_{\rm S}$ between $10^{-4}-10^{-1}$, then angular momentum loss through GW r-mode instability will dominate the spin-down of the pulsar. Whereas, extremely large deformation of $\epsilon\gtrsim 10^{-3}$ is required for the spin-down of a typical millisecond pulsar to be dominated by GW emission due to time-dependent mass quadrupolar moment, which has been ruled out under various physical conditions. Our analysis shows that if the $\epsilon\lesssim 10^{-7}$ the spin-down is dominated by the magnetic dipole radiation in the absence of r-mode instability, which is consistent with the current limit of the GW observations.
@@
We extended our study in the light of millisecond magnetars, which are believed to be born from NS-NS mergers, core-collapse of a massive star or accretion-induced collapse of a white dwarf. The spin-down energy is believed to power the X-ray plateau phase of GRBs. These millisecond magnetars can have large magnetic field strength $\sim 10^{15}$ G. We find at large spin-frequencies $\gtrsim 600$ Hz the presence of large amplitude r-mode oscillations dominates the spin-down rate. However, the effect of rapid rotation suppresses the braking index closer to 6 instead of 7. Whereas, the impact of non-axisymmetric deformation on braking index is less pronounced unless $\epsilon>10^{-3}$. However, if the birth spin frequency is low $\sim 200$ Hz or r-mode amplitude is less than $10^{-1}$, spin-down is always dominated by the magnetic dipole radiation.
@@
We have used the constrained EOS to compute the permissible range of the braking index as a function of the rotational frequency. The width of the range of the braking index increases with rotation frequency, indicating a wide range of EOS-dependent structural change of neutron stars in response to large centrifugal force. We find a millisecond magnetar with the magnetic field strength $10^{15}$ G, deformation $\epsilon=10^{-7}$ and r-mode amplitude $\alpha_{\rm S}=10^{-4}$ has braking index $\sim 3$ at low frequency and $n<3$ at higher frequency [^4] It is important to emphasise that millisecond magnetar are prone to magnetospheric changes Contopoulos and Spitkovsky (2006), where the separatrix radius (distance from centre to the edge of the closed magnetosphere, beyond which the magnetic field lines open) can lag behind the radius of light cylinder (distance from the centre to the radius of the rigid co-rotating magnetosphere), which naturally lead to $n<3$ scenario. However, this modification in the magnetosphere is beyond the scope of this paper and has been confined to the simple dipolar approximation of the magnetosphere. as shown in Fig. 7. Our band of computed braking is in good agreement with the measured braking index of GRB 140903A and 130603B, from where we find the spin frequency of the central engine of GRB 140903A could be between $\sim 550-850$ Hz, which is subjected to systematics at higher spin-frequency end. Our model of spin-frequency dependent braking index hints towards a potential method to constrain the dense mater EOS by precisely measuring the braking index and the spin frequency from the high-energy GRB light curves.
@@ Acknowledgements
Pulsar research at Jodrell Bank Centre for Astrophysics and Jodrell Bank Observatory is supported by a consolidated grant from the UK Science and Technology Facilities Council (STFC). P.C. is supported by European Union’s HORIZON MSCA-2022-PF-01-01 Programme under Grant Agreement No. 101109652, project ProMatEx-NS. P.C. also acknowledges past support of the Fonds de la Recherche Scientifique-FNRS, Belgium, under Grant No. 4.4501.19. Part of the computation has been performed in the high-performance computing facility “Magus,” which is available at the Shiv Nadar Institution of Eminence.

### v2

@@ V Conclusion
In this paper, we have investigated the effect of stellar rotation on the braking index and its dependence on the EOS and baryon mass. We have generated a large number of EOSs by employing a semi-agnostic model and constrained them using the latest astrophysical observations from the gravitational and radio wave bands within a Bayesian framework. Our constrained parameter space is also consistent with the observational findings of neutron star’s mass and radius from X-ray bands. We then use these constrained EOS to study the effect on the macroscopic quantities of rapidly rotating neutron stars like mass, radius, moment of inertia and the ratio of polar to equatorial radius. Our analysis shows neutron stars with stiffer EOS experience larger deformations in the shape at a given frequency, in comparison to the softer EOS. Consequently, neutron stars with stiffer EOS have smaller Kepler frequencies.
@@
Next, we study the effect of rotation on the braking index of neutron stars. We present the first calculation of braking index in the presence of the r-mode instability jointly with the other braking mechanism like MDR and GWs due to non-axisymmetric deformations, and incorporating the rotational effect. We have shown that the rotation of the star can impact its braking index significantly. For example, the rotational effects make the braking index smaller than 3 in the case of magnetic dipole radiation when the spin frequency $\gtrsim 200$ Hz. However, the low braking index in normal pulsars has been explained through various other mechanisms, such as, the evolution of magnetic inclination angle, which has not been considered here. Similarly, we have shown that if the neutron stars spin-down is purely due to GW radiation from non-asymmetric deformations or r-mode oscillations, rapid rotation leads to the significant departure of braking index from their classical values of 5 and 7, respectively.
@@
We further explore the dependence of the braking index on the rotation frequency for various physically motivated combinations of surface magnetic field strength, ellipticity of the star, and the amplitude of r-mode oscillations. We find if there is r-mode instability generated in a typical millisecond pulsar with $\alpha_{\rm S}$ between $10^{-4}-10^{-1}$, then angular momentum loss through GW r-mode instability will dominate the spin-down of the pulsar. Whereas, extremely large deformation of $\epsilon\gtrsim 10^{-3}$ is required for the spin-down of a typical millisecond pulsar to be dominated by GW emission due to time-dependent mass quadrupolar moment, which has been ruled out under various physical conditions. Our analysis shows that if the $\epsilon\lesssim 10^{-7}$, the spin-down is dominated by the magnetic dipole radiation in the absence of r-mode instability, which is consistent with the current limit of the GW observations.
@@
We extended our study in the light of millisecond magnetars, which are believed to be born from NS-NS mergers, core-collapse of a massive star, or accretion-induced collapse of a white dwarf. The spin-down energy is believed to power the X-ray plateau phase of GRBs. These millisecond magnetars can have large magnetic field strength $\sim 10^{15}$ G. We find that at large spin-frequencies $\gtrsim 600$ Hz, the presence of large amplitude r-mode oscillations dominates the spin-down rate. However, the effect of rapid rotation suppresses the braking index closer to 6 instead of 7. But the impact of non-axisymmetric deformation on the braking index is less pronounced unless $\epsilon>10^{-3}$. However, if the birth spin frequency is low $\sim 200$ Hz or r-mode amplitude is less than $10^{-1}$, spin-down is always dominated by the magnetic dipole radiation.
@@
We have used the constrained EOS to compute the permissible range of the braking index as a function of the rotational frequency. The width of the range of the braking index increases with rotation frequency, indicating a wide range of EOS-dependent structural change of neutron stars in response to large centrifugal force. We find a millisecond magnetar with the magnetic field strength $10^{15}$ G, deformation $\epsilon=10^{-7}$ and r-mode amplitude $\alpha_{\rm S}=10^{-4}$ has braking index $\sim 3$ at low frequency and $n<3$ at higher frequency [^4] It is important to emphasize that millisecond magnetar are prone to magnetospheric changes Contopoulos and Spitkovsky (2006), where the separatrix radius (distance from center to the edge of the closed magnetosphere, beyond which the magnetic field lines open) can lag behind the radius of light cylinder (distance from the center to the radius of the rigid co-rotating magnetosphere), which naturally lead to $n<3$ scenario. However, this modification in the magnetosphere is beyond the scope of this paper and has been confined to the simple dipolar approximation of the magnetosphere. as shown in Fig. 7. We have further investigated the effect of baryon mass on the spin-frequency dependent braking index and found that massive stars have a larger braking index compared to the star with lower baryon mass at a given spin-frequency. We find there exists a degeneracy between the uncertainty due to choice of baryon mass and the uncertainty in the EOS. This work provides a framework for future rigorous analysis for modeling GRB light curves.
@@ Acknowledgements
Pulsar research at Jodrell Bank Centre for Astrophysics and Jodrell Bank Observatory is supported by a consolidated grant from the UK Science and Technology Facilities Council (STFC). This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101034371. PC acknowledges the support from the European Union’s HORIZON MSCA-2022-PF-01-01 Programme under Grant Agreement No. 101109652, project ProMatEx-NS. Part of the computation has been performed in the high-performance computing facility “Magus,” which is available at the Shiv Nadar Institution of Eminence.